In each of the following cases, calculate the accounting break-even and the cash break-even points. Ignore any tax effects in calculating the cash break-even. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)In each of the following cases, calculate the accounting break-even and the cash break- even points. Ignore any tax effects i
Answer #1

Accounting Break even = Fixed Costs including depreciation / Contribution p.u.

Cash Break even = Fixed Costs excluding Depreciation / Contribution p.u

Contribution p.u. = Sales p.u. - Variable Cost p.u.

Case 1,

Accounting Breakeven = (8,120,000 + 3,060,000) / (3,370 - 2,675) = 16,086.33 units

Cash Breakeven = 8,120,000 / 695 = 11,683.45 units

CASE 2,

Accounting Breakeven = (78,000 + 350,000) / (146 - 81)

= 6,584.62 units

Cash Breakeven = 78,000 / 65 = 1,200 units

Case 3

Accounting Breakeven = (3700 + 860) / (31 - 7)

= 190 units

Cash Breakeven = 3700 / 24 = 154.17 units